### 20 Jan 2024 v1

# Construction of the log-convex minorant of a sequence $\{M_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$

Abstract.
We give a simple construction of the log-convex minorant of a sequence $\{M_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ and consequently extend to the $d$-dimensional case the well-known formula that relates a log-convex sequence $\{M_{p}\}_{p\in\mathbb{N}_{0}}$ to its associated function $\omega_{M}$, that is $M_{p}=\sup_{t>0}t^{p}\exp(-\omega_{M}(t))$. We show that in the more dimensional anisotropic case the classical log-convex condition $M_{\alpha}^{2}\leq M_{\alpha-e_{j}}M_{\alpha+e_{j}}$ is not sufficient: convexity as a function of more variables is needed (not only coordinate-wise). We finally obtain some applications to the inclusion of spaces of rapidly decreasing ultradifferentiable functions in the matrix weighted setting.
Key words and phrases:
regularization of sequences, log-convex sequences, ultradifferentiable functions, matrix weights
2020 Mathematics Subject Classification:
40B05, 46A13, 46A45
January 20, 2024 14:48
1. Introduction
For a sequence $\{M_{p}\}_{p\in\mathbb{N}_{0}}$ of real positive numbers (with $M_{0}=1$ for simplicity; $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$), its associated function is defined by
$\omega_{M}(t):=\sup_{p\in\mathbb{N}_{0}}\log\frac{t^{p}}{M_{p}},\qquad t>0.$
Mandelbrojt proved in [M, Chap. I] (see also [K]) that if $\lim_{p\to+\infty}M_{p}^{1/p}=+\infty$ then
(1.1) $M_{p}=\sup_{t>0}\frac{t^{p}}{\exp\omega_{M}(t)},\qquad p\in\mathbb{N}_{0},$
if and only if $\{M_{p}\}_{p\in\mathbb{N}_{0}}$ is logarithmically convex, i.e.
$M_{p}^{2}\leq M_{p-1}M_{p+1},\qquad\forall p\in\mathbb{N}.$
We refer also to the recent work [S] where this construction has been studied again in detail, some technical ambiguities have been solved and non-standard cases have been studied as well; see Section 3.
However, to the best of our knowledge this condition has never been generalized to the $d$-dimensional anisotropic case ($d>1$), and the reason is that the classical coordinate-wise logarithmic convexity condition (5.8) for a sequence $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ is not sufficient to obtain the analogous of (1.1) for $M_{\alpha}$, as explained in Remark 5.3 (here $e_{j}$ is the $j$-th element of the canonical basis of $\mathbb{R}^{d}$ with all entries equal to 0 except the $j$-th entry equal to 1). The reason is that this is a convexity condition on each variable separately and not on the globality of its variables. Assuming the stronger condition that $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ is log-convex on the globality of its variables $\alpha\in\mathbb{N}_{0}^{d}$ (see Definition 5.1), we extend (1.1) to $\alpha\in\mathbb{N}_{0}^{d}$ instead of $p\in\mathbb{N}_{0}$ (see Theorem 5.2).
To obtain this result we construct in Sections 2-4 the (optimal) convex minorant of a sequence $\{a_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ (then $a_{\alpha}=\log M_{\alpha}$ in Section 5). The idea, in the one variable case, takes inspiration from the convex regularization of sequences of Mandelbrojt in [M], which was quite complicated and difficult to export to the more dimensional case. Our construction is made by taking the supremum of hyperplanes approaching from below the given sequence and leads to the notion of convexity for a sequence $\{a_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ in the sense that $a_{\alpha}=F(\alpha)$ for a convex function $F:[0,+\infty)^{d}\to\mathbb{R}$. This condition gives the suitable notion of logarithmic convexity for a sequence $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ in order to write it in terms of its associated function as in (5.7).
As an application of the above results, in Section 6 we characterize the condition on the matrix weights $\mathcal{M}$ and $\mathcal{N}$ in order to have a continuous inclusion between the spaces $\mathcal{S}_{\{\mathcal{M}\}}/\mathcal{S}_{(\mathcal{M})}$, $\mathcal{S}_{\{\mathcal{N}\}}/\mathcal{S}_{(\mathcal{N})}$ of rapidly decreasing ultradifferentiable functions in the matrix weighted setting, both in the Roumieu and Beurling cases.
2. Construction of the convex minorant candidate
Let us recall that a real sequence $\{a_{p}\}_{p\in\mathbb{N}_{0}}$ is said to be convex if
$a_{p}\leq\frac{1}{2}a_{p-1}+\frac{1}{2}a_{p+1},\qquad\forall p\in\mathbb{N}.$
This is equivalent to say that the polygonal obtained by connecting the points $(p,a_{p})$ with $(p+1,a_{p+1})$ by a straight line, for all $p\in\mathbb{N}_{0}$, is the graph of a convex function, or equivalently that there exists a convex function $F:\,[0,+\infty)\to\mathbb{R}$ with $F(p)=a_{p}$ for all $p\in\mathbb{N}_{0}$.
This suggests for a real sequence $\{a_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ the following:
Definition 2.1.
We say that a sequence $\{a_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ is convex if there exists a convex function $F:\,[0,+\infty)^{d}\to\mathbb{R}$ with $F(\alpha)=a_{\alpha}$ for all $\alpha\in\mathbb{N}_{0}^{d}$.
Definition 2.2.
The convex minorant of a sequence $\{a_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ is the largest convex sequence $\{a^{c}_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ with $a^{c}_{\alpha}\leq a_{\alpha}$ for all $\alpha\in\mathbb{N}_{0}^{d}$.
We want to construct the convex minorant of a sequence $\{a_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}\subset\overline{\mathbb{R}}:=\mathbb{R}\cup\{\pm\infty\}$ such that:
  1. $(i)$
    $a_{\alpha}>-\infty,\quad\forall\alpha\in\mathbb{N}_{0}^{d},$
  2. $(ii)$
    $\lim_{|\alpha|\to+\infty}\frac{a_{\alpha}}{|\alpha|}=+\infty,\quad\mbox{for}\ |\alpha|=\alpha_{1}+\ldots+\alpha_{d},$
  3. $(iii)$
    $a_{\alpha}$ may be +$\infty$+ ∞ at most for a finite number of multi-indices $\alpha\in\mathbb{N}_{0}^{d}$,
  4. $(iv)$
    $a_{0}\in\mathbb{R}.$
To this aim we set
$S:=$$\{(\alpha,a_{\alpha}):\ \alpha\in\mathbb{N}_{0}^{d}\},$
$\mathcal{L}:=$$\{f:\mathbb{R}^{d}\to\mathbb{R}:\ f\ \mbox{is an affine function}\}$
$=$ $\{f(x)=\langle k,x\rangle+c:\ (k,c)\in\mathbb{R}^{d}\times\mathbb{R}\}.$
Note that the graphs of the affine functions $f\in\mathcal{L}$ are hyperplanes in $\mathbb{R}^{d+1}$.
Lemma 2.3.
Given $f\in\mathcal{L}$ we have
$f(\alpha)>a_{\alpha},$
at most for a finite number of points $\alpha\in\mathbb{N}_{0}^{d}$.
Proof.
For $|\alpha|\geq 1$ by the Cauchy-Schwarz inequality we have
$\frac{|f(\alpha)|}{|\alpha|}\leq\frac{|\langle k,\alpha\rangle|+|c|}{|\alpha|}\leq\frac{\|k\|\cdot\|\alpha\|}{|\alpha|}+|c|\leq\|k\|+|c|,$
since $\|\alpha\|=\sqrt{\alpha_{1}^{2}+\ldots+\alpha_{d}^{2}}\leq\alpha_{1}+\ldots+\alpha_{d}=|\alpha|$.
On the other hand, assumption $(ii)$ implies
$\frac{a_{\alpha}}{|\alpha|}>\|k\|+|c|,$
for $|\alpha|$ large enough. This implies that only a finite number of $\alpha\in\mathbb{N}_{0}^{d}$ may satisfy that $|f(\alpha)|>a_{\alpha}$. ∎
Let us now consider
$\mathcal{L}_{S}:=\{f\in\mathcal{L}:\ f(\alpha)\leq a_{\alpha}\ \forall\alpha\in\mathbb{N}_{0}^{d}\}.$
The graphs of the functions $f\in\mathcal{L}_{S}$ are the hyperplanes which lie under $S$.
Note that $\mathcal{L}_{S}\neq\emptyset$ by Lemma 2.3. As a matter of fact, given $f\in\mathcal{L}$, since $f(\alpha)\leq a_{\alpha}$ except a finite number of points $\alpha_{1},\ldots,\alpha_{\ell}$, we have that
$f-\max_{1\leq j\leq\ell}\{f(\alpha_{j})-a_{\alpha_{j}}\}\in\mathcal{L}_{S}.$
The idea is now to consider the supremum of these hyperplanes which lie under $S$ and project on this set each $a_{\alpha}$ to construct the desired convex minorant sequence. So let us first define
(2.1) $F(x):=\sup_{f\in\mathcal{L}_{S}}f(x).$
Note that the above defined $F:\mathbb{R}^{d}\to\mathbb{R}$ is a convex function since it is the supremum of a set of convex (affine) functions: the epigraph is convex because it is the intersection of convex sets. By construction
$F(\alpha)\leq a_{\alpha},\qquad\forall\alpha\in\mathbb{N}_{0}^{d},$
and we claim that
$F(x)<+\infty\qquad\forall x\in[0,+\infty)^{d}.$
As a matter of fact, given $x\in[0,+\infty)^{d}$, by assumptions $(i)$ and $(iii)$, we can find $d+1$ points $\alpha_{1},\ldots,\alpha_{d+1}\in\mathbb{N}_{0}^{d}$ such that $x$ is inside (or on the border of) the simplex of vertices $\alpha_{1},\ldots,\alpha_{d+1}$ and $f(\alpha_{j})\leq a_{\alpha_{j}}<+\infty$ for every $f\in\mathcal{L}_{S}$ and $11\leq j\leq d+$ 1. Then
$f(x)\leq\max_{1\leq j\leq d+1}f(\alpha_{j})<+\infty,$
by the convexity of $f$, and hence $F(x)<+\infty$.
Note also that $F$ is continuous on $(0,+\infty)^{d}$, being a convex function.
For fixed $k\in\mathbb{R}^{d}$ let us now define
(2.2) $h_{k}:=$ $\sup\{c\in\mathbb{R}:\ f(x)=\langle k,x\rangle+c\in\mathcal{L}_{S}\}$
$=$ $\sup\{c\in\mathbb{R}:\ \langle k,\alpha\rangle+c\leq a_{\alpha}\,,\ \forall\alpha\in\mathbb{N}_{0}^{d}\}.$
We have:
Lemma 2.4.
Let $k\in\mathbb{R}^{d}$ and $h_{k}$ as in (2.2). Then
(2.3) $h_{k}=\inf_{\alpha\in\mathbb{N}_{0}^{d}}\{a_{\alpha}-\langle k,\alpha\rangle\}=\min_{\alpha\in\mathbb{N}_{0}^{d}}\{a_{\alpha}-\langle k,\alpha\rangle\}.$
Proof.
Let us first remark that $\inf_{\alpha\in\mathbb{N}_{0}^{d}}\{a_{\alpha}-\langle k,\alpha\rangle\}$ is a minimum because assumption $(ii)$ implies that
$\lim_{|\alpha|\to+\infty}(a_{\alpha}-\langle k,\alpha\rangle)=+\infty,$
(see the proof of Lemma 2.3), so that the infimum is attained on a bounded set of $\mathbb{N}_{0}^{d}$, which is a finite set.
Let us now set
(2.4) $\tilde{h}_{k}:=\inf_{\alpha\in\mathbb{N}_{0}^{d}}\{a_{\alpha}-\langle k,\alpha\rangle\}=\min_{\alpha\in\mathbb{N}_{0}^{d}}\{a_{\alpha}-\langle k,\alpha\rangle\}=a_{\bar{\alpha}}-\langle k,\bar{\alpha}\rangle$
for some minimum point $\bar{\alpha}\in\mathbb{N}_{0}^{d}$, and prove that $\tilde{h}_{k}=h_{k}$ as defined in (2.2).
Clearly
$\tilde{h}_{k}\leq a_{\alpha}-\langle k,\alpha\rangle,\qquad\forall\alpha\in\mathbb{N}_{0}^{d},$
[…]
Therefore
$f(x)=\tilde{h}_{k}+\langle k,x\rangle\in\mathcal{L}_{S}$
is one of the functions in (2.2) with $c=\tilde{h}_{k}$ and hence $h_{k}\geq\tilde{h}_{k}$.
On the contrary, if $c>\tilde{h}_{k}$ then (2.4) implies that
$c+\langle k,\bar{\alpha}\rangle>\tilde{h}_{k}+\langle k,\bar{\alpha}\rangle=a_{\bar{\alpha}}$
so that in this case
$\bar{f}(x)=c+\langle k,x\rangle\notin\mathcal{L}_{S}$
and therefore $h_{k}\leq\tilde{h}_{k}$. ∎
Now we define
(2.5) $f_{k}(x):=h_{k}+\langle k,x\rangle.$
Proposition 2.5.
Let $F$ be as in (2.1), $h_{k}$ as in (2.3) and $f_{k}$ as in (2.5). Then
$F(x)=\sup_{k\in\mathbb{R}^{d}}f_{k}(x).$
Proof.
Since $f_{k}\in\mathcal{L}_{S}$ by (2.3), we clearly have that
$F(x)\geq\sup_{k\in\mathbb{R}^{d}}f_{k}(x).$
On the other hand, if $f\in\mathcal{L}_{S}$ then
$f(x)=c+\langle k,x\rangle$
for some $c\in\mathbb{R}$, $k\in\mathbb{R}^{d}$, and hence by (2.2)
$f(x)\leq f_{k}(x)\leq\sup_{\ell\in\mathbb{R}^{d}}f_{\ell}(x).$
This finally implies that
$F(x)=\sup_{f\in\mathcal{L}_{S}}f(x)\leq\sup_{k\in\mathbb{R}^{d}}f_{k}(x).$
We thus have a convex sequence $\{\tilde{a}_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ defined, for $F$ as in (2.1), by
(2.6) $\tilde{a}_{\alpha}:=F(\alpha),$
and which can be equivalently be defined by
$\tilde{a}_{\alpha}:=\sup_{k\in\mathbb{R}^{d}}(\langle k,\alpha\rangle+h_{k}),$
for
$h_{k}=\inf_{\alpha\in\mathbb{N}_{0}^{d}}(a_{\alpha}-\langle k,\alpha\rangle).$
We call it the convex minorant candidate of $\{a_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ since we shall prove in the following sections that it is indeed the convex minorant of $\{a_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ (see Corollary 4.2).
The first step in this direction is to clarify the one-dimensional case (in Section 3) and then proceed by descending induction on $d$ for the $d$-dimensional case (in Section 4). To this aim in the next section we shall look more closely at the geometric construction in the one dimensional case, that is particularly clear.
3. Geometric construction of the convex minorant sequence in the one-dimensional case
The geometric construction in the one-dimensional case can also be found in [S]; we refer to this work and in particular to [S, Sect. 3.3] concerning the comparison with the classical results from [M, Chap. I]. Here we revisit the construction in the spirit of Section 2 since it is required in the induction argument for the higher-dimensional situation in Section 4.
Let $\{a_{\alpha}\}_{\alpha\in\mathbb{N}_{0}}$ satisfy $(i)-(iv)$. In particular the condition $a_{0}\in\mathbb{R}$ will be essential for the first step of the construction. All functions of $\mathcal{L}_{S}$ are of the form
$f(x)=c+kx,\qquad\mbox{with}\ k\in\mathbb{R}\ \mbox{and}\ c\leq a_{0},$
since $f(\alpha)\leq a_{\alpha}$ for all $\alpha\in\mathbb{N}_{0}$ implies, in particular, $f(0)=c\leq a_{0}$.
Let us now consider the functions of $\mathcal{L}_{S}$ of the form
$f_{a_{0},k}(x)=a_{0}+kx,\qquad k\in\mathbb{R},$
and note that
$F(0)=\sup_{f\in\mathcal{L}_{S}}f(0)=a_{0}.$
The idea is now to rotate (i.e. increasing the slope $k$) this straight line $y=a_{0}+kx$ around the point $(0,a_{0})\in S$ until we meet another point $(p,a_{p})\in S$. In order to have $f_{a_{0},k}(p)=a_{p}$ we find
$a_{p}=a_{0}+kp\quad\Leftrightarrow\quad k=\frac{a_{p}-a_{0}}{p}.$
Take then
$k_{0}:=\inf_{p\in\mathbb{N}}\frac{a_{p}-a_{0}}{p}$
and note that it is a minimum because of the assumption $a_{p}/p\to+\infty$ for $p\to+\infty$, which implies that the infimum can be done on a bounded subset of $\mathbb{N}$ and hence on a finite number of indices $p$:
$k_{0}=\min_{p\in\mathbb{N}}\frac{a_{p}-a_{0}}{p}=\frac{a_{p_{1}}-a_{0}}{p_{1}}$
for some $p_{1}\in\mathbb{N}$. Observe that $p_{1}$ does not need to be unique; if there is more than one $p_{1}$ realizing the minimum, for the construction below it doesn’t matter which one we choose at this step.
Set
(3.1) $f_{a_{0},k_{0}}(x):=a_{0}+k_{0}x.$
We claim that
$F(x)=f_{a_{0},k_{0}}(x)=a_{0}+k_{0}x,\qquad\forall x\in[0,p_{1}],$
where $F$ is the function defined in (2.1), i.e. the geometric construction coincides with the construction made in §2, in $[0,p_{1}]$.
On the one side, $f_{a_{0},k_{0}}\in\mathcal{L}_{S}$ by construction and hence
$F(x)\geq f_{a_{0},k_{0}}(x),\qquad\forall x\in[0,p_{1}].$
On the other side, if
$f(x)=c+kx\in\mathcal{L}_{S}$
then we must have
$f(0)\leq a_{0}=f_{a_{0},k_{0}}(0)$
$f(p_{1})\leq a_{p_{1}}=f_{a_{0},k_{0}}(p_{1})$
and hence
$f(x)\leq f_{a_{0},k_{0}}(x),\qquad\forall x\in[0,p_{1}],$
since they are affine functions. Therefore
$F(x)=\sup_{f\in\mathcal{L}_{S}}f(x)\leq f_{a_{0},k_{0}}(x),\qquad\forall x\in[0,p_{1}].$
We have thus proved that
$F(x)=a_{0}+k_{0}x,\qquad\forall x\in[0,p_{1}],$
and moreover
$F(p_{1})=a_{p_{1}}.$
We can further proceed as in the previous step, considering
$f_{a_{p_{1}},k}(x)=a_{p_{1}}+k(x-p_{1})$
with $f_{a_{p_{1}},k}(p_{1})=a_{p_{1}}=F(p_{1})$. Requiring
$f_{a_{p_{1}},k}(p)=a_{p}\ \Leftrightarrow\ a_{p}=a_{p_{1}}+k(p-p_{1})\ \Leftrightarrow\ k=\frac{a_{p}-a_{p_{1}}}{p-p_{1}}$
we take
$k_{1}:=\inf_{p_{1}<p\in\mathbb{N}}\frac{a_{p}-a_{p_{1}}}{p-p_{1}}=\min_{p_{1}<p\in\mathbb{N}}\frac{a_{p}-a_{p_{1}}}{p-p_{1}}=\frac{a_{p_{2}}-a_{p_{1}}}{p_{2}-p_{1}},$
for some $p_{1}<p_{2}\in\mathbb{N}$, and set
$f_{a_{p_{1}},k_{1}}(x):=a_{p_{1}}+k_{1}(x-p_{1}).$
Then, for all $p\in[p_{1},p_{2}]$,
$F(x)=$ $f_{a_{p_{1}},k_{1}}(x)=a_{p_{1}}+k_{1}(x-p_{1})$
$=$ $k_{1}x+a_{p_{1}}-k_{1}p_{1}$
$=$ $k_{1}x+a_{p_{1}}-\frac{a_{p_{2}}-a_{p_{1}}}{p_{2}-p_{1}}p_{1}$
$=$ $k_{1}x+\frac{p_{2}a_{p_{1}}-p_{1}a_{p_{2}}}{p_{2}-p_{1}}$
$=$ $k_{1}x+d_{k_{1}}$
with
$d_{k_{1}}=\frac{p_{2}a_{p_{1}}-p_{1}a_{p_{2}}}{p_{2}-p_{1}}.$
Moreover $F(p_{2})=a_{p_{2}}$. Also in this case, $p_{2}$ does not need to be unique, and the choice of $p_{2}$ does not affect the next steps.
Going on recursively in the same way we have a geometric construction which coincides with the construction of $F$ in §2.
The convex minorant candidate sequence given by $\tilde{a}_{p}=F(p)$ as defined in (2.6) is in this case the projection of $a_{p}$ on the segments of lines above defined in each interval $[p_{i},p_{i+1}]$:
$\tilde{a}_{p_{i}}=$ $a_{p_{i}},\qquad\forall i\in\mathbb{N}_{0},$
$\tilde{a}_{p}=$ $pk_{i}+d_{k_{i}}$
$=$ $\frac{a_{p_{i+1}}-a_{p_{i}}}{p_{i+1}-p_{i}}\,p+\frac{p_{i+1}a_{p_{i}}-p_{i}a_{p_{i+1}}}{p_{i+1}-p_{i}},\qquad p_{i}<p<p_{i+1}.$
In the one dimensional case it is thus immediate by the construction that the convex minorant candidate sequence $\{\tilde{a}_{p}\}_{p\in\mathbb{N}_{0}}$ coincides with the convex minorant $\{a^{c}_{p}\}_{p\in\mathbb{N}_{0}}$.
4. Convex minorant sequence in the ${d}$-dimensional case
In order to show that the convex minorant candidate sequence $\{\tilde{a}_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}:=\{F(\alpha)\}_{\alpha\in\mathbb{N}_{0}^{d}}$ defined in (2.6) is the convex minorant sequence of $\{a_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$, we shall now prove that $F$ is the biggest convex function whose epigraph contains $S$.
Theorem 4.1.
The function $F$ defined in (2.1) coincides with the biggest convex function $g:[0,+\infty)^{d}\to\mathbb{R}$ such that
(4.1) $g(\alpha)\leq a_{\alpha},\qquad\forall\alpha\in\mathbb{N}_{0}^{d}.$
Proof.
Let us first remark that, since $F$ is a convex function such that
$F(\alpha)\leq a_{\alpha},\qquad\forall\alpha\in\mathbb{N}_{0}^{d},$
the largest convex function $g$ which satisfies (4.1) must be larger than $F$:
(4.2) $g(x)\geq F(x),\qquad\forall x\in[0,+\infty)^{d}.$
In order to prove the opposite inequality, let us first work in the interior of $[0,+\infty)^{d}$. Fix $x^{0}\in(0,+\infty)^{d}$ and consider $(x^{0},y^{0})=(x^{0},g(x^{0}))$ on the graph of $g$.
Since $g$ is convex, its epigraph $G_{g}^{+}$ is a convex subset of $\mathbb{R}^{d+1}$. It follows, as a consequence of the Hahn-Banach theorem, that there is a hyperplane of the form
$y=\langle k^{*},x-x^{0}\rangle+y^{0}=\langle k^{*},x\rangle+c^{*},$
for some $k^{*}\in\mathbb{R}^{d}$ and $c^{*}=y^{0}-\langle k^{*},x^{0}\rangle$, that leaves the whole set $G_{g}^{+}$ on the same side of the hyperplane. Note that $x^{0}\in(0,+\infty)^{d}$ avoids “vertical” hyperplanes of the form $x_{j}=c$, because of the convexity of $g$.
Then
$f^{*}(x):=\langle k^{*},x\rangle+c^{*}\in\mathcal{L}_{S},$
since $g(\alpha)\leq a_{\alpha}$ for all $\alpha\in\mathbb{N}_{0}^{d}$ by assumption. It follows that
$F(x^{0})=\sup_{f\in\mathcal{L}_{S}}f(x^{0})\geq f^{*}(x^{0})=y^{0}=g(x^{0}),$
and hence, by the arbitrariness of $x^{0}\in(0,+\infty)^{d}$ and by (4.2), we have that
(4.3) $F(x^{0})=g(x^{0}),\qquad\forall x^{0}\in(0,+\infty)^{d}.$
Let us now consider the case $x^{0}\in\partial[0,+\infty)^{d}$, i.e. $x^{0}_{j}=0$ for at least one $1\leq j\leq d$. Assume for simplicity $j=d$ and set $x=(x^{\prime},0)=(x_{1},\ldots,x_{d-1},0)$ for $x\in[0,+\infty)^{d-1}\times\{0\}$.
The problem in this case is that the graph of $g$ could have in $(x^{0},g(x^{0}))$ a tangent “vertical” hyperplane of the form $\{x_{d}=0\}$ and hence not defined by a function of $\mathcal{L}_{S}$. The idea is to prove that the trace of $F$ on $\{x_{d}=0\}$ is a supremum of affine functions on $\mathbb{R}^{d-1}$ whose graph is below $S\cap\{x_{d}=0\}$ reducing to the case of dimension $d-1$, and then proceed recursively up to dimension 1, where the assumption that $a_{0}\in\mathbb{R}$ guarantees the conclusion.
So let us define
$\mathcal{L}_{S}^{0}:=\{f(x^{\prime})=\langle k^{\prime},x^{\prime}\rangle+c:\ k^{\prime}\in\mathbb{R}^{d-1},c\in\mathbb{R},\langle k^{\prime},\alpha^{\prime}\rangle+c\leq a_{(\alpha^{\prime},0)}\ \forall\alpha^{\prime}\in\mathbb{N}_{0}^{d-1}\}.$
We claim that
(4.4) $F(x^{\prime},0)=\sup_{f\in\mathcal{L}_{S}^{0}}f(x^{\prime})=:F^{0}(x^{\prime}).$
In the following, we use the convention that when $k\in\mathbb{R}^{d}$ we put $k^{\prime}\in\mathbb{R}^{d-1}$ with $k=(k^{\prime},k_{d})$. Note first that Proposition 2.5 for the sequence $\{a_{(\alpha^{\prime},0)}\}_{\alpha^{\prime}\in\mathbb{N}_{0}^{d-1}}$ implies that
$\sup_{f\in\mathcal{L}_{S}^{0}}f(x^{\prime})=\sup_{k^{\prime}\in\mathbb{R}^{d-1}}(\langle k^{\prime},x^{\prime}\rangle+h^{0}_{k^{\prime}})$
with
$h^{0}_{k^{\prime}}=\min_{\alpha^{\prime}\in\mathbb{N}_{0}^{d-1}}(a_{(\alpha^{\prime},0)}-\langle k^{\prime},\alpha^{\prime}\rangle).$
Moreover, by Proposition 2.5 once more,
$F(x^{\prime},0)=\sup_{f\in\mathcal{L}_{S}}f(x^{\prime},0)=\sup_{k\in\mathbb{R}^{d}}(\langle k^{\prime},x^{\prime}\rangle+h_{k})$
with
$h_{k}=\min_{\alpha\in\mathbb{N}_{0}^{d}}(a_{\alpha}-\langle k,\alpha\rangle)\leq\min_{(\alpha^{\prime},0)\in\mathbb{N}_{0}^{d-1}\times\{0\}}(a_{(\alpha^{\prime},0)}-\langle k,(\alpha^{\prime},0)\rangle)=h^{0}_{k^{\prime}}$
and hence
$F(x^{\prime},0)\leq$ $\sup_{k\in\mathbb{R}^{d}}(\langle k^{\prime},x^{\prime}\rangle+h^{0}_{k^{\prime}})=\sup_{k^{\prime}\in\mathbb{R}^{d-1}}(\langle k^{\prime},x^{\prime}\rangle+h^{0}_{k^{\prime}})$
$=$ $\sup_{f\in\mathcal{L}_{S}^{0}}f(x^{\prime})=F^{0}(x^{\prime}).$
In order to prove the opposite inequality we remark that if $f\in\mathcal{L}_{S}$, i.e.
$f(x)=\langle k,x\rangle+c,\quad k\in\mathbb{R}^{d},\,c\in\mathbb{R},$
(4.5) with $\langle k,\alpha\rangle+c\leq a_{\alpha},\quad\forall\alpha\in\mathbb{N}_{0}^{d},$
then
$f(x^{\prime},0)=\langle k^{\prime},x^{\prime}\rangle+c\in\mathcal{L}_{S}^{0}$
since $\langle k^{\prime},\alpha^{\prime}\rangle+c\leq a_{(\alpha^{\prime},0)}$ by (4.5).
Therefore
$F^{0}(x^{\prime})=\sup_{h\in\mathcal{L}_{S}^{0}}h(x^{\prime})\geq\sup_{f\in\mathcal{L}_{S}}f(x^{\prime},0)=F(x^{\prime},0)$
and the equality (4.4) is proved. This means that we have reduced the problem to prove that $F^{0}(x^{\prime})$ is the maximum convex function $g:[0,+\infty)^{d-1}\to\mathbb{R}$ such that
$g(\alpha^{\prime})\leq a_{(\alpha^{\prime},0)}\qquad\forall\alpha^{\prime}\in\mathbb{N}_{0}^{d-1}.$
If $x^{0}=(x_{1}^{0},\ldots,x_{d-1}^{0})\in(0,+\infty)^{d-1}$ the thesis follows from (4.3) (in the case of dimension $d-1$ instead of $d$). If $x^{0}\in\partial[0,+\infty)^{d-1}$, for instance $x^{0}=(x_{1}^{0},\ldots,x_{d-2}^{0},0)$ we argue as before and thus reduce to determine the biggest convex function $g:[0,+\infty)^{d-2}\to\mathbb{R}$ with
$g(\alpha_{1},\ldots,\alpha_{d-2})\leq a_{(\alpha_{1},\ldots,\alpha_{d-2},0,0)},\ \forall(\alpha_{1},\ldots,\alpha_{d-2})\in\mathbb{N}_{0}^{d-2}.$
Proceeding recursively we are finally lead to the one-dimensional case for the sequence $\{a_{(\alpha_{1},0,\ldots,0)}\}_{\alpha_{1}\in\mathbb{N}_{0}}$, where the construction of §3 gives the desired maximum convex function whose graph is below $S$. Note that on $[0,+\infty)$ the problem of the border does not appear since $a_{0}\in\mathbb{R}$ by assumption $(iv)$ and the first line through $(0,a_{0})$ is not a vertical line, but the graph of
$f_{a_{0},k_{o}}(x)=a_{0}+k_{0}x$
defined in (3.1). In particular $F(0)=f_{a_{0},k_{0}}(0)=a_{0}$. ∎
Corollary 4.2.
Given a sequence $\{a_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ satisfying $(i)-(iv)$, its convex minorant sequence $\{a^{c}_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ is defined by
$a^{c}_{\alpha}=F(\alpha),$
for $F$ as in (2.1), or equivalently by
(4.6) $a^{c}_{\alpha}:=\sup_{k\in\mathbb{R}^{d}}(\langle k,\alpha\rangle+h_{k}),$
for
(4.7) $h_{k}=\inf_{\alpha\in\mathbb{N}_{0}^{d}}(a_{\alpha}-\langle k,\alpha\rangle).$
In particular $a_{0}^{c}=a_{0}$.
5. Construction of the log-convex minorant
Let $\{M_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ be a sequence of positive real numbers such that
(5.1) $\lim_{|\alpha|\to+\infty}M_{\alpha}^{1/|\alpha|}=+\infty$
(we can also allow $M_{\alpha}=+\infty$ for finitely many multi-indices $\alpha\neq 0$). We say that the sequence $\mathbf{M}=\{M_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ is normalized if $M_{0}=1$.
For a normalized sequence $\mathbf{M}=\{M_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ we define its associated function $\omega_{\mathbf{M}}$ by
$\omega_{\mathbf{M}}(t):=\sup_{\alpha\in\mathbb{N}_{0,t}^{d}}\log\frac{|t^{\alpha}|}{M_{\alpha}},\qquad t\in\mathbb{R}^{d},$
with
$\mathbb{N}_{0,t}^{d}:=\{\alpha\in\mathbb{N}_{0}^{d}:\ \alpha_{j}=0\ \mbox{if}\ t_{j}=0,\,j=1,\ldots,d\},$
and the convention that $10^{0}=$ 1. We observe that here the supremum is made on $\mathbb{N}_{0,t}^{d}$ in order to ensure that in the definition of the associated function the argument of the logarithm is not 00; we can equivalently write
$\omega_{\mathbf{M}}(t)=\sup_{\alpha\in\mathbb{N}^{d}_{0}}\log\frac{|t^{\alpha}|}{M_{\alpha}},\qquad t\in\mathbb{R}^{d},$
with the convention that $\log 0=-\infty$. In what follows (in particular in Section 6) we use the latter expression for convenience.
Condition (5.1) ensures that $\omega_{\mathbf{M}}(t)<+\infty$ for all $t\in\mathbb{R}^{d}$ (see [M, Chap. I] or [BJOS1, Rem. 1]).
The function $\omega_{\mathbf{M}}(t)$ is increasing on $(0,+\infty)^{d}$ in the following sense: $\omega_{\mathbf{M}}(t)\leq\omega_{\mathbf{M}}(s)$ if $t\leq s$ with the order relation $t_{j}\leq s_{j}$ for all $1\leq j\leq d$.
Consider then
(5.2) $a_{\alpha}=\log M_{\alpha},\qquad\alpha\in\mathbb{N}_{0}^{d}.$
By (5.1) the sequence $\{a_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ satisfies all assumptions $(i)-(iv)$ of §2.
We can thus consider the convex minorant sequence $\{a_{\alpha}^{c}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ as in (4.6)-(4.7) and call
$M^{\mathop{\rm lc}\nolimits}_{\alpha}:=\exp a^{c}_{\alpha},\qquad\forall\alpha\in\mathbb{N}_{0}^{d},$
the log-convex minorant of $\{M_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$.
By the results of the previous section $\{\log M^{\mathop{\rm lc}\nolimits}_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ is the largest convex sequence less than or equal to $\{a_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ defined in (5.2) and they coincide if $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ is log-convex, according to the following
Definition 5.1.
A sequence $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ is said to be log-convex if there exists a convex function $F:[0,+\infty)^{d}\to\mathbb{R}$ with $F(\alpha)=\log M_{\alpha}$ for all $\alpha\in\mathbb{N}_{0}^{d}$.
By construction
(5.3) $a^{c}_{\alpha}=\sup_{k\in\mathbb{R}^{d}}(\langle k,\alpha\rangle+h_{k})=\sup_{k\in\mathbb{R}^{d}}(\langle k,\alpha\rangle-A(k)),$
where the so-called trace function $A(k)$ is given by
$A(k)=-h_{k}=$ $-\inf_{\alpha\in\mathbb{N}_{0}^{d}}(a_{\alpha}-\langle k,\alpha\rangle)$
$=$ $\sup_{\alpha\in\mathbb{N}_{0}^{d}}(\langle k,\alpha\rangle-a_{\alpha})$
$=$ $\sup_{\alpha\in\mathbb{N}_{0}^{d}}(\langle k,\alpha\rangle-\log M_{\alpha})$
$=$ $\sup_{\alpha\in\mathbb{N}_{0}^{d}}\log\frac{e^{\langle k,\alpha\rangle}}{M_{\alpha}}$
$=$ $\sup_{\alpha\in\mathbb{N}_{0}^{d}}\log\frac{|(e^{k})^{\alpha}|}{M_{\alpha}}$
(5.4) $=$ $\omega_{\mathbf{M}}(e^{k}),$
since
$|(e^{k})^{\alpha}|=|(e^{k_{1}},\ldots,e^{k_{d}})^{\alpha}|=|e^{k_{1}\alpha_{1}}\cdots e^{k_{d}\alpha_{d}}|=e^{\langle k,\alpha\rangle}.$
From (5.3) and (5.4) we have that
$M^{\mathop{\rm lc}\nolimits}_{\alpha}=$ $\exp a^{c}_{\alpha}=\exp\{\sup_{k\in\mathbb{R}^{d}}(\langle k,\alpha\rangle-\omega_{\mathbf{M}}(e^{k}))\}$
$=$ $\sup_{k\in\mathbb{R}^{d}}\frac{|(e^{k})^{\alpha}|}{\exp\omega_{\mathbf{M}}(e^{k})}$
$=$ $\sup_{s\in(0,+\infty)^{d}}\frac{|s^{\alpha}|}{\exp\omega_{\mathbf{M}}(s)}$
(5.5) $=$ $\sup_{s\in(0,+\infty)^{d}}\frac{s^{\alpha}}{\exp\omega_{\mathbf{M}}(s)}.$
In particular, being $M_{\alpha}^{\mathop{\rm lc}\nolimits}\leq M_{\alpha}$ for any normalized sequence of positive real numbers $\{M_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$, we have
(5.6) $\sup_{s\in(0,+\infty)^{d}}\frac{s^{\alpha}}{\exp\omega_{\mathbf{M}}(s)}\leq M_{\alpha},\qquad\forall\alpha\in\mathbb{N}_{0}^{d},$
and, moreover, $\{M_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ is log-convex if and only if the equality holds. Hence, we have proved:
Theorem 5.2.
Let $\{M_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ be a normalized sequence of positive real numbers satisfying (5.1). Then, the sequence $\{M_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ is log-convex if and only if
(5.7) $M_{\alpha}=\sup_{s\in(0,+\infty)^{d}}\frac{s^{\alpha}}{\exp\omega_{\mathbf{M}}(s)},\qquad\forall\alpha\in\mathbb{N}_{0}^{d}.$
Remark 5.3.
Note that if $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ is a log-convex sequence, i.e. $\log M_{\alpha}=F(\alpha)$ for a convex function $F:[0,+\infty)^{d}\to\mathbb{R}$, then
$F(\alpha)=F\left(\frac{1}{2}(\alpha-e_{j})+\frac{1}{2}(\alpha+e_{j})\right)\leq\frac{1}{2}F(\alpha-e_{j})+\frac{1}{2}F(\alpha+e_{j}),$
that is
$2\log M_{\alpha}\leq\log M_{\alpha-e_{j}}+\log M_{\alpha+e_{j}}.$
This yields the coordinate-wise log-convexity condition for a sequence $\{M_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ :
[…]
This condition is clearly weaker than the condition of logarithmic convexity given in Definition 5.1 since there are functions which are coordinate-wise convex but not convex as functions of more variables. In particular, from (5.5) we have that (5.8) is not sufficient to obtain (5.7); for an explicit example, see Example 5.4 below. Clearly in the one-dimensional case the two notions of log-convexity coincide, and (5.7) was already known (see [K, Prop. 3.2]).
Example 5.4.
The function of two variables $F(x,y)=(x+1)^{2}(y+1)^{2}$ is coordinate-wise convex but not convex in $[0,+\infty)^{2}$, and the sequence defined by $M_{\alpha}:=e^{F(\alpha)-1}$ is normalized and satisfies (5.1), but does not satisfy (5.7).
6. Characterization of inclusion relations of spaces of rapidly decreasing ultradifferentiable functions
Let us first recall the notion of weight matrices (anisotropic framework) as considered in [BJOS1, Sec. 3].
A weight matrix $\mathcal{M}$ is the set
:=$\mathcal{M}:=\{(\mathbf{M}^{(\lambda)})_{\lambda>0}:\mathbf{M}^{(\lambda)}=(M_{\alpha}^{(\lambda)})_{\alpha\in\mathbb{N}_{0}^{d}},M_{0}^{(\lambda)}=1,M_{\alpha}^{(\lambda)}\leq M_{\alpha}^{(\kappa)}\ \forall\alpha\in\mathbb{N}_{0}^{d}\ \forall 0<\lambda\leq\kappa\}.$
Denoting by $\|\cdot\|_{\infty}$ the supremum norm we consider the following spaces of matrix weighted global ultradifferentiable functions of Roumieu/Beurling type
$\mathcal{S}_{\{\mathcal{M}\}}(\mathbb{R}^{d}):=\{f\in C^{\infty}(\mathbb{R}^{d}):\ \exists\lambda,h,C>0:\ \sup_{\alpha,\beta\in\mathbb{N}_{0}^{d}}\frac{\|x^{\alpha}\partial^{\beta}f\|_{\infty}}{h^{|\alpha+\beta|}M_{\alpha+\beta}^{(\lambda)}}\leq C\}$
$\mathcal{S}_{(\mathcal{M})}(\mathbb{R}^{d}):=\{f\in C^{\infty}(\mathbb{R}^{d}):\ \forall\lambda,h>0\,\exists C_{h,\lambda}>0:\ \sup_{\alpha,\beta\in\mathbb{N}_{0}^{d}}\frac{\|x^{\alpha}\partial^{\beta}f\|_{\infty}}{h^{|\alpha+\beta|}M_{\alpha+\beta}^{(\lambda)}}\leq C_{h,\lambda}\}$
endowed with the inductive limit topology in the Roumieu case and the projective limit topology in the Beurling case (see [BJOS1, Sec. 3]).
In order to characterize the inclusion of spaces of this type, given two weight matrices =$\mathcal{M}=\{(\mathbf{M}^{(\lambda)})_{\lambda>0}\}$ and $\mathcal{N}=\{(\mathbf{N}^{(\lambda)})_{\lambda>0}\}$ we introduce the following relations:
{}$\mathcal{M}\{\preceq\}\mathcal{N}\quad\mbox{if}\quad\forall\lambda>0\exists\kappa>0\exists C\geq 1\,\mbox{s.t.}\ M_{\alpha}^{(\lambda)}\leq C^{|\alpha|}N_{\alpha}^{(\kappa)}\ \forall\alpha\in\mathbb{N}_{0}^{d}$
$\mathcal{M}(\preceq)\mathcal{N}\quad\mbox{if}\quad\forall\lambda>0\exists\kappa>0\exists C\geq 1\,\mbox{s.t.}\ M_{\alpha}^{(\kappa)}\leq C^{|\alpha|}N_{\alpha}^{(\lambda)}\ \forall\alpha\in\mathbb{N}_{0}^{d}$
$\mathcal{M}\vartriangleleft\mathcal{N}\quad\mbox{if}\quad\forall\lambda>0\forall\kappa>0\forall h>0\exists C\geq 1\,\mbox{s.t.}\ M_{\alpha}^{(\lambda)}\leq Ch^{|\alpha|}N_{\alpha}^{(\kappa)}\ \forall\alpha\in\mathbb{N}_{0}^{d}.$
We shall also need the following conditions for a weight matrix =$\mathcal{M}=\{(\mathbf{M}^{(\lambda)})_{\lambda>0}\}$ (see [BJOS1]; cf. also [L]), in the Roumieu setting
(6.1) $\forall\lambda>0\exists\kappa\geq\lambda,B,C,H>0:\quad\alpha^{\alpha/2}M_{\beta}^{(\lambda)}\leq BC^{|\alpha|}H^{|\alpha+\beta|}M_{\alpha+\beta}^{(\kappa)}\ \ \forall\alpha,\beta\in\mathbb{N}_{0}^{d}$
(6.2) $\forall\lambda>0\exists\kappa\geq\lambda,A\geq 1:\quad M_{\alpha+e_{j}}^{(\lambda)}\leq A^{|\alpha|+1}M_{\alpha}^{(\kappa)}\ \ \forall\alpha\in\mathbb{N}_{0}^{d},1\leq j\leq d$
(6.3) $\forall\lambda>0\exists\kappa\geq\lambda,A\geq 1:\quad M_{\alpha}^{(\lambda)}M_{\beta}^{(\lambda)}\leq A^{|\alpha+\beta|}M_{\alpha+\beta}^{(\kappa)}\ \ \forall\alpha,\beta\in\mathbb{N}_{0}^{d}$
and in the Beurling case
$\forall\lambda>0\exists 0<\kappa\leq\lambda,H>0:\forall C>0\exists B>0:$
(6.4) $\hskip 56.9055pt\alpha^{\alpha/2}M_{\beta}^{(\kappa)}\leq BC^{|\alpha|}H^{|\alpha+\beta|}M_{\alpha+\beta}^{(\lambda)}\ \ \forall\alpha,\beta\in\mathbb{N}_{0}^{d}$
$\forall\lambda>0\exists 0<\kappa\leq\lambda,A\geq 1:$
(6.5) $\hskip 56.9055ptM_{\alpha+e_{j}}^{(\kappa)}\leq A^{|\alpha|+1}M_{\alpha}^{(\lambda)}\ \ \forall\alpha\in\mathbb{N}_{0}^{d},1\leq j\leq d$
$\forall\lambda>0\exists 0<\kappa\leq\lambda,A\geq 1:$
(6.6) $\hskip 56.9055ptM_{\alpha}^{(\kappa)}M_{\beta}^{(\kappa)}\leq A^{|\alpha+\beta|}M_{\alpha+\beta}^{(\lambda)}\ \ \forall\alpha,\beta\in\mathbb{N}_{0}^{d}.$
Note that (6.1) for $\beta=0$ implies that $\mathbf{M}^{(\kappa)}$ satisfies (5.1) for some $\kappa>0$, and hence for all $\kappa^{\prime}\geq\kappa$. Similarly (6.4) implies that $\mathbf{M}^{(\lambda)}$ satisfies (5.1) for all $\lambda>0$ (see [BJOS1, Rem. 3]). Since condition (5.1) ensures that the associate weight function $\omega_{\mathbf{M}}$ is finite, the above remarks are essential to recall, from [BJOS1, Thm. 1], that if the weight matrix =$\mathcal{M}=\{(\mathbf{M}^{(\lambda)})_{\lambda>0}\}$ satisfies (6.1) and (6.2) then $\mathcal{S}_{\{\mathcal{M}\}}$ is isomorphic to the sequence space
(6.7) $\Lambda_{\{\mathcal{M}\}}:=\{\mathbf{c}=(c_{\alpha})_{\alpha\in\mathbb{N}_{0}^{d}}\in\mathbb{C}^{\mathbb{N}_{0}^{d}}:\ \exists\lambda,h>0\,\mbox{s.t.}\,\sup_{\alpha\in\mathbb{N}_{0}^{d}}|c_{\alpha}|e^{\omega_{\mathbf{M}^{(\lambda)}}(\alpha^{1/2}/h)}<+\infty\},$
and similarly in the Beurling case, for a weight matrix $\mathcal{M}$ having (6.4)-(6.5), the space $\mathcal{S}_{(\mathcal{M})}$ is isomorphic to
(6.8) $\Lambda_{(\mathcal{M})}:=\{\mathbf{c}=(c_{\alpha})_{\alpha\in\mathbb{N}_{0}^{d}}\in\mathbb{C}^{\mathbb{N}_{0}^{d}}:\ \forall\lambda,h>0\,\sup_{\alpha\in\mathbb{N}_{0}^{d}}|c_{\alpha}|e^{\omega_{\mathbf{M}^{(\lambda)}}(\alpha^{1/2}/h)}<+\infty\},$
where $\alpha^{1/2}:=(\alpha_{1}^{1/2},\ldots,\alpha_{d}^{1/2})$.
It will also be useful, in the following, to write the sequence spaces (6.7)-(6.8) as follows (see [BJOS1]):
(6.10) $\Lambda_{\{\mathcal{M}\}}=\{\mathbf{c}=(c_{\alpha})_{\alpha\in\mathbb{N}_{0}^{d}}\in\mathbb{C}^{\mathbb{N}_{0}^{d}}:\ \exists j\in\mathbb{N}\,\mbox{s.t.}\,\sup_{\alpha\in\mathbb{N}_{0}^{d}}|c_{\alpha}|e^{\omega_{\mathbf{M}^{(j)}}(\alpha^{1/2}/j)}<+\infty\}$
$\Lambda_{(\mathcal{M})}=\{\mathbf{c}=(c_{\alpha})_{\alpha\in\mathbb{N}_{0}^{d}}\in\mathbb{C}^{\mathbb{N}_{0}^{d}}:\ \forall j\in\mathbb{N}\,\sup_{\alpha\in\mathbb{N}_{0}^{d}}|c_{\alpha}|e^{\omega_{\mathbf{M}^{(1/j)}}(j\alpha^{1/2})}<+\infty\}.$
We say that a weight matrix =$\mathcal{M}=\{(\mathbf{M}^{(\lambda)})_{\lambda>0}\}$ is log-convex if $\{M^{(\lambda)}_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$ is log-convex (according to Definition 5.1) for all $\lambda>0$.
Let us remark that in the one-dimensional case the assumption of log-convexity, together with $M_{0}=1$, implies both (6.3) and (6.6) with $A=1$ (we are here considering the sequence case for simplicity, i.e., the case when in the weight matrix $\mathcal{M}$ all the $\mathbf{M}^{(\lambda)}$, $\lambda>0$, coincide) since the convex function $F(p)=\log M_{p}$ has increasing difference quotient and hence
$\frac{\log M_{p+q}-\log M_{p}}{q}\geq\frac{\log M_{q}-\log M_{0}}{q}=\frac{\log M_{q}}{q}.$
On the contrary, in the more-dimensional case log-convexity and $M_{0}=1$ do not imply (6.3)/(6.6), not even under the additional conditions (6.1)/(6.4) and (6.2)/(6.5). Take, for instance, $M_{\alpha}=\alpha^{\alpha/2}e^{\max\{\alpha_{1}^{2},\alpha_{2}^{2}\}}$ for $\alpha=(\alpha_{1},\alpha_{2})\in\mathbb{N}_{0}^{2}$ (with the convention $10^{0}:=$ 1). It is easy to check that it is log-convex (since $\log M_{\alpha}$ is the sum of convex functions) and satisfies (6.1) and (6.2). However, taking $\alpha=(n,0)$ and $\beta=(0,n)$ condition (6.3) is not valid for $n\to+\infty$.
The above remarks explain why conditions (6.3) and (6.6), that we shall need in Theorem 6.1, are not required in the one-dimensional/isotropic case (see [BJOS2]).
Theorem 6.1.
Let =$\mathcal{M}=\{(\mathbf{M}^{(\lambda)})_{\lambda>0}\}$ and $\mathcal{N}=\{(\mathbf{N}^{(\lambda)})_{\lambda>0}\}$ be two weight matrices and assume that $\mathcal{M}$ is log-convex. Then the following characterizations hold:
  1. (i)
    Let $\mathcal{M}$ satisfy (6.1)-(6.3) and $\mathcal{N}$ satisfy (6.1)-(6.2). Then: $\mathcal{S}_{\{\mathcal{M}\}}(\mathbb{R}^{d})\subseteq\mathcal{S}_{\{\mathcal{N}\}}(\mathbb{R}^{d})$ holds with continuous inclusion if and only if {}$\mathcal{M}\{\preceq\}\mathcal{N}$;
  2. (ii)
    Let $\mathcal{M}$ satisfy (6.1)-(6.3) and $\mathcal{N}$ satisfy (6.4)-(6.5). Then: $\mathcal{S}_{\{\mathcal{M}\}}(\mathbb{R}^{d})\subseteq\mathcal{S}_{(\mathcal{N})}(\mathbb{R}^{d})$ holds with continuous inclusion if and only if $\mathcal{M}\vartriangleleft\mathcal{N}$;
  3. (iii)
    Let $\mathcal{M}$ satisfy (6.4)-(6.6) and $\mathcal{N}$ satisfy (6.4)-(6.5). Then: $\mathcal{S}_{(\mathcal{M})}(\mathbb{R}^{d})\subseteq\mathcal{S}_{(\mathcal{N})}(\mathbb{R}^{d})$ holds with continuous inclusion if and only if $\mathcal{M}(\preceq)\mathcal{N}$.
Proof.
The sufficiency of condition {}$\mathcal{M}\{\preceq\}\mathcal{N}$ (resp. $\mathcal{M}\vartriangleleft\mathcal{N}$, $\mathcal{M}(\preceq)\mathcal{N}$) for the inclusion of the spaces clearly follows from the definition of the spaces (hypotheses (6.1)-(6.3)/(6.4)-(6.6) and log-convexity of $\mathcal{M}$ are not needed at this step).
We now prove the necessity of the conditions.
$(i)$: Let us assume that $\mathcal{S}_{\{\mathcal{M}\}}(\mathbb{R}^{d})\subseteq\mathcal{S}_{\{\mathcal{N}\}}(\mathbb{R}^{d})$ holds with continuous inclusion and prove that {}$\mathcal{M}\{\preceq\}\mathcal{N}$.
By the already mentioned isomorphism with the sequence space (6.10), we have that
(6.11) $\Lambda_{\{\mathcal{M}\}}\simeq\mathcal{S}_{\{\mathcal{M}\}}(\mathbb{R}^{d})\subseteq\mathcal{S}_{\{\mathcal{N}\}}(\mathbb{R}^{d})\simeq\Lambda_{\{\mathcal{N}\}}.$
For some fixed (but arbitrary) $j\in\mathbb{N}$, $j\geq 2$, let us consider the sequence $\mathbf{c}:=(c_{\alpha})_{\alpha\in\mathbb{N}_{0}^{d}}\in\mathbb{C}^{\mathbb{N}_{0}^{d}}$ defined by
(6.12) $c_{\alpha}:=e^{-\omega_{\mathbf{M}^{(j)}}(\bar{\alpha}^{1/2}/j)}$
for $\bar{\alpha}=(\bar{\alpha}_{1},\ldots,\bar{\alpha}_{d})$ with
(6.13) $\bar{\alpha}_{k}=\begin{cases}\alpha_{k}&\mbox{if}\ \alpha_{k}\neq 1\cr 0&\mbox{if}\ \alpha_{k}=1,\end{cases}$
with the convention $e^{-\infty}=0$.
Let us prove that $\mathbf{c}\in\Lambda_{\{\mathcal{M}\}}$. If $\omega_{\mathbf{M}^{(j)}}(\bar{\alpha}^{1/2}/j)=+\infty$ there is nothing to prove. If $\omega_{\mathbf{M}^{(j)}}(\bar{\alpha}^{1/2}/j)<+\infty$ we must find $h\in\mathbb{N}$ and $C>0$ such that
$|c_{\alpha}|=e^{-\omega_{\mathbf{M}^{(j)}}(\bar{\alpha}^{1/2}/j)}\leq Ce^{-\omega_{\mathbf{M}^{(h)}}(\alpha^{1/2}/h)}\qquad\forall\alpha\in\mathbb{N}_{0}^{d},$
i.e.
(6.14) $\omega_{\mathbf{M}^{(h)}}\left(\frac{\sqrt{\alpha_{1}}}{h},\ldots,\frac{\sqrt{\alpha_{d}}}{h}\right)\leq\log C+\omega_{\mathbf{M}^{(j)}}\left(\frac{\sqrt{\bar{\alpha}_{1}}}{j},\ldots,\frac{\sqrt{\bar{\alpha}_{d}}}{j}\right).$
Let us first choose $\bar{h}\in\mathbb{N}$, $\bar{h}\geq j$, such that $\mathbf{M}^{(h)}$ satisfies (5.1), and hence $\omega_{\mathbf{M}^{(h)}}(t)<+\infty$ for all $t\in\mathbb{R}^{d}$, for all $h\geq\bar{h}$ (this is possible because of assumption (6.1)). Then by assumption (6.3) we choose $h\in\mathbb{N}$, $h\geq\bar{h}$, and $A\geq 1$ such that
(6.15) $M^{(j)}_{\alpha}M^{(j)}_{\beta}\leq A^{|\alpha+\beta|}M^{(h)}_{\alpha+\beta},\qquad\forall\alpha,\beta\in\mathbb{N}_{0}^{d}.$
For such a choice of $h$ and from the definition of associated function
(6.16) $\omega_{\mathbf{M}^{(h)}}\left(\frac{\sqrt{\alpha_{1}}}{h},\ldots,\frac{\sqrt{\alpha_{d}}}{h}\right)=\sup_{\beta_{1},\ldots,\beta_{d}\in\mathbb{N}_{0}}\log\frac{\alpha_{1}^{\beta_{1}/2}\cdots\alpha_{d}^{\beta_{d}/2}}{h^{\beta_{1}+\ldots+\beta_{d}}M^{(h)}_{(\beta_{1},\ldots,\beta_{d})}}.$
Let us now assume, without loss of generality, that the entries of $\alpha$ equal to 1 (if there are some) are in the first positions, i.e. $(\alpha_{1},\ldots,\alpha_{d})=(1,\ldots,1,\alpha_{s+1},\ldots,\alpha_{d})$, for some $s\geq 0$, with $\alpha_{k}\neq 1$ for $s+1\leq k\leq d$. By (6.15) we have that
$M^{(h)}_{(\beta_{1},\ldots,\beta_{d})}\geq\frac{1}{A^{\beta_{1}+\ldots+\beta_{d}}}M^{(j)}_{(\beta_{1},\ldots,\beta_{s},0,\ldots,0)}M^{(j)}_{(0,\ldots,0,\beta_{s+1},\ldots,\beta_{d})}$
and hence from (6.16)
$\omega_{\mathbf{M}^{(h)}}\left(\frac{\alpha^{1/2}}{h}\right)\leq$ $\sup_{\beta_{1},\ldots,\beta_{d}\in\mathbb{N}_{0}}\Bigg{(}\log\frac{\alpha_{s+1}^{\beta_{s+1}/2}\cdots\alpha_{d}^{\beta_{d}/2}}{\left(\frac{h}{A}\right)^{\beta_{s+1}+\ldots+\beta_{d}}M^{(j)}_{(0,\ldots,0,\beta_{s+1},\ldots,\beta_{d})}}$
$\qquad+\log\frac{(A/h)^{\beta_{1}+\ldots+\beta_{s}}}{M^{(j)}_{(\beta_{1},\ldots,\beta_{s},0,\ldots,0)}}\Bigg{)}$
$\leq$ $\sup_{\beta_{s+1},\ldots,\beta_{d}\in\mathbb{N}_{0}}\log\frac{\alpha_{s+1}^{\beta_{s+1}/2}\cdots\alpha_{d}^{\beta_{d}/2}}{\left(\frac{h}{A}\right)^{\beta_{s+1}+\ldots+\beta_{d}}M^{(j)}_{(0,\ldots,0,\beta_{s+1},\ldots,\beta_{d})}}$
$+\sup_{\beta_{1},\ldots,\beta_{s}\in\mathbb{N}_{0}}\log\frac{\left(\frac{A}{h}\right)^{\beta_{1}+\ldots+\beta_{s}}}{M^{(j)}_{(\beta_{1},\ldots,\beta_{s},0,\ldots,0)}}.$
Eventually enlarging $h$ so that $h\geq Aj$ we thus have
$\omega_{\mathbf{M}^{(h)}}\left(\frac{\alpha^{1/2}}{h}\right)\leq$ $\sup_{\beta_{s+1},\ldots,\beta_{d}\in\mathbb{N}_{0}}\log\frac{\alpha_{s+1}^{\beta_{s+1}/2}\cdots\alpha_{d}^{\beta_{d}/2}}{j^{\beta_{s+1}+\ldots+\beta_{d}}M^{(j)}_{(0,\ldots,0,\beta_{s+1},\ldots,\beta_{d})}}+\omega_{\bar{\mathbf{M}}^{(j)}}\left(\frac{A}{h}\right)$
(6.17) $\leq$ $\omega_{\mathbf{M}^{(j)}}\left(\frac{\bar{\alpha}^{1/2}}{j}\right)+C_{j,h}$
for $\{\bar{M}_{\beta}^{(j)}\}_{\beta\in\mathbb{N}_{0}^{s}}$ the sequence given by $\bar{M}_{\beta}^{(j)}=M^{(j)}_{(\beta_{1},\ldots,\beta_{s},0,\ldots,0)}$ and $C_{j,h}=\omega_{\bar{\mathbf{M}}^{(j)}}(A/h)$. Note that this constant is finite because if $|(\beta_{1},\ldots,\beta_{s})|\to+\infty$ then also $|(\beta_{1},\ldots,\beta_{s},0,\ldots,0)|\to+\infty$ and hence
$\left(\bar{M}^{j}_{\beta}\right)^{1/|\beta|}=\left({M}^{j}_{(\beta_{1},\ldots,\beta_{s},0,\ldots,0)}\right)^{1/(\beta_{1}+\cdots+\beta_{s})}\longrightarrow+\infty$
by (5.1).
Inequality (6.17) proves (6.14) and hence $\mathbf{c}\in\Lambda_{\{\mathcal{M}\}}$.
From (6.11) we have that $\mathbf{c}\in\Lambda_{\{\mathcal{N}\}}$, and therefore there exist $\ell\in\mathbb{N}$ and $C\geq 1$ such that
$e^{-\omega_{\mathbf{M}^{(j)}}(\bar{\alpha}^{1/2}/j)}=|c_{\alpha}|\leq Ce^{-\omega_{\mathbf{N}^{(\ell)}}(\alpha^{1/2}/\ell)},\qquad\forall\alpha\in\mathbb{N}_{0}^{d},$
i.e.
(6.18) $\omega_{\mathbf{N}^{(\ell)}}(\alpha^{1/2}/\ell)\leq\log C+\omega_{\mathbf{M}^{(j)}}(\bar{\alpha}^{1/2}/j),\qquad\forall\alpha\in\mathbb{N}_{0}^{d}.$
Fix now $t\in(0,+\infty)^{d}$ and set $\mathbf{1}:=(1,\ldots,1)\in\mathbb{R}^{d}$. There exists $\alpha\in\mathbb{N}_{0}^{d}$ with $\alpha^{1/2}<t\leq(\alpha+\mathbf{1})^{1/2}$, i.e. $\alpha_{k}<t_{k}^{2}\leq\alpha_{k}+1$ for all $1\leq k\leq d$. From (6.18) we have
$\omega_{\mathbf{N}^{(\ell)}}(t/\ell)\leq$ $\omega_{\mathbf{N}^{(\ell)}}((\alpha+\mathbf{1})^{1/2}/\ell)\leq\log C+\omega_{\mathbf{M}^{(j)}}((\overline{\alpha+\mathbf{1}})^{1/2}/j)$
$\leq$ $\log C+\omega_{\mathbf{M}^{(j)}}(\alpha^{1/2})\leq\log C+\omega_{\mathbf{M}^{(j)}}(t)$
because $\omega_{\mathbf{M}^{(j)}}$ is increasing on $(0,+\infty)^{d}$ and $(\overline{\alpha+\mathbf{1}})^{1/2}/j\leq\alpha^{1/2}$ since $\overline{\alpha_{k}+1}=0$ if $\alpha_{k}=0$ and $(\overline{\alpha_{k}+1})^{1/2}/j=\sqrt{\alpha_{k}+1}/j\leq\sqrt{\alpha_{k}}$ if $\alpha_{k}\geq 1$ and $j\geq 2$.
From (5.6) and (5.7) we thus have that
$M^{(j)}_{\alpha}=$ $\sup_{t\in(0,+\infty)^{d}}\frac{t^{\alpha}}{\exp\omega_{\mathbf{M}^{(j)}}(t)}\leq C\sup_{t\in(0,+\infty)^{d}}\frac{t^{\alpha}}{\exp\omega_{\mathbf{N}^{(\ell)}}(t/\ell)}$
$=$ $C\sup_{s\in(0,+\infty)^{d}}\frac{(s\ell)^{\alpha}}{\exp\omega_{\mathbf{N}^{(\ell)}}(s)}\leq C\ell^{|\alpha|}N_{\alpha}^{(\ell)}.$
We have thus proved that
$\forall j\in\mathbb{N},j\geq 2,\exists\ell\in\mathbb{N},C\geq 1:\quad M^{(j)}_{\alpha}\leq C\ell^{|\alpha|}N_{\alpha}^{(\ell)}\quad\forall\alpha\in\mathbb{N}_{0}^{d}.$
Since $M_{\alpha}^{(j)}\leq M_{\alpha}^{(2)}$ for $j<2$ and the sequences in the weight matrices are normalized, we have proved that {}$\mathcal{M}\{\preceq\}\mathcal{N}$.
$(ii)$: Assuming
$\Lambda_{\{\mathcal{M}\}}\simeq\mathcal{S}_{\{\mathcal{M}\}}(\mathbb{R}^{d})\subseteq\mathcal{S}_{(\mathcal{N})}(\mathbb{R}^{d})\simeq\Lambda_{(\mathcal{N})}$
we have to prove that $\mathcal{M}\vartriangleleft\mathcal{N}$.
We choose $\mathbf{c}=(c_{\alpha})_{\alpha\in\mathbb{N}_{0}^{d}}\in\mathbb{C}^{\mathbb{N}^{d}}$ as in (6.12) so that $\mathbf{c}\in\Lambda_{\{\mathcal{M}\}}\subseteq\Lambda_{(\mathcal{N})}(\mathbb{R}^{d})$ i.e.
(6.19) $\qquad\ \forall j\in\mathbb{N},j\geq 2,\forall\ell\in\mathbb{N}\,\exists C\geq 1:\ \ e^{-\omega_{\mathbf{M}^{(j)}}(\bar{\alpha}^{1/2}/j)}=|c_{\alpha}|\leq Ce^{-\omega_{\mathbf{N}^{(1/\ell)}}(\alpha^{1/2}\ell)}\ \forall\alpha\in\mathbb{N}_{0}^{d}.$
As in the case $(i)$, for any $t\in(0,+\infty)^{d}$ we can choose $\alpha\in\mathbb{N}_{0}^{d}$ with $\alpha^{1/2}<t\leq(\alpha+\mathbf{1})^{1/2}$ so that from (6.19):
$\omega_{\mathbf{N}^{(1/\ell)}}(t\ell)\leq$ $\omega_{\mathbf{N}^{(1/\ell)}}((\alpha+\mathbf{1})^{1/2}\ell)\leq\log C+\omega_{\mathbf{M}^{(j)}}((\overline{\alpha+\mathbf{1}})^{1/2}/j)$
$\leq$ $\log C+\omega_{\mathbf{M}^{(j)}}(\alpha^{1/2})\leq\log C+\omega_{\mathbf{M}^{(j)}}(t).$
It follows that for all $j\in\mathbb{N}$, $j\geq 2$, and for all $\ell\in\mathbb{N}$ there exists $C\geq 1$ such that
$M^{(j)}_{\alpha}=$ $\sup_{t\in(0,+\infty)^{d}}\frac{t^{\alpha}}{\exp\omega_{\mathbf{M}^{(j)}}(t)}\leq C\sup_{t\in(0,+\infty)^{d}}\frac{t^{\alpha}}{\exp\omega_{\mathbf{N}^{(1/\ell)}}(t\ell)}$
$=$ $C\sup_{s\in(0,+\infty)^{d}}\frac{(s/\ell)^{\alpha}}{\exp\omega_{\mathbf{N}^{(1/\ell)}}(s)}\leq C\frac{1}{\ell^{|\alpha|}}N_{\alpha}^{(1/\ell)},$
because of (5.6) and (5.7). Since $M_{\alpha}^{(j)}\leq M_{\alpha}^{(2)}$ for $j<2$ we finally obtain
(6.20) $\forall j\in\mathbb{N}\,\forall\ell\in\mathbb{N}\,\exists C\geq 1:\quad M_{\alpha}^{(j)}\leq C\frac{1}{\ell^{|\alpha|}}N_{\alpha}^{(1/\ell)}\ \ \forall\alpha\in\mathbb{N}_{0}^{d}.$
Now, it is obvious that (6.20) implies condition $\mathcal{M}\vartriangleleft\mathcal{N}$. In fact, it is enough to take $\ell\geq\max\{1/\kappa,1/h\}$ for any given $\kappa,h>0$ in the definition of $\mathcal{M}\vartriangleleft\mathcal{N}$.
$(iii)$: Assuming
$\Lambda_{(\mathcal{M})}\simeq\mathcal{S}_{(\mathcal{M})}(\mathbb{R}^{d})\subseteq\mathcal{S}_{(\mathcal{N})}(\mathbb{R}^{d})\simeq\Lambda_{(\mathcal{N})}$
we have to prove that $\mathcal{M}(\preceq)\mathcal{N}$.
By the continuity of the inclusion and (6.10) we have
(6.21) $\begin{split}&\forall\ell\in\mathbb{N}\,\exists h\in\mathbb{N},C\geq 1\,\mbox{s.t.}\,\forall\mathbf{c}\in\Lambda_{(\mathcal{M})}\\ &\sup_{\alpha\in\mathbb{N}_{0}^{d}}|c_{\alpha}|e^{\omega_{\mathbf{N}^{(1/\ell)}}(\ell\alpha^{1/2})}\leq C\sup_{\alpha\in\mathbb{N}_{0}^{d}}|c_{\alpha}|e^{\omega_{\mathbf{M}^{(1/h)}}(h\alpha^{1/2})}.\end{split}$
Let us now fix $\alpha\in\mathbb{N}_{0}^{d}$ and assume as before, without loss of generality, that $(\alpha_{1},\ldots,\alpha_{d})=(1,\ldots,1,\alpha_{s+1},\ldots,\alpha_{d})$, for some $s\geq 0$, with $\alpha_{k}\neq 1$ for $s+1\leq k\leq d$. Setting $\bar{\alpha}$ as in (6.13), by (6.6) there exist $j\in\mathbb{N}$, $j\geq h$, and $A\geq 1$ such that
$\omega_{\mathbf{M}^{(1/h)}}(h\alpha^{1/2})=$ $\sup_{\beta\in\mathbb{N}_{0}^{d}}\log\frac{h^{\beta_{1}+\ldots+\beta_{d}}\alpha_{1}^{\beta_{1}/2}\cdots\alpha_{d}^{\beta_{d}/2}}{M^{(1/h)}_{\beta}}$
$\leq$ $\sup_{\beta\in\mathbb{N}_{0}^{d}}\log\frac{h^{\beta_{1}+\ldots+\beta_{d}}\cdot 1\cdot\alpha_{s+1}^{\beta_{s+1}/2}\cdots\alpha_{d}^{\beta_{d}/2}}{(1/A)^{\beta_{1}+\ldots+\beta_{d}}M^{(1/j)}_{(\beta_{1},\ldots,\beta_{s},0,\ldots 0)}M^{(1/j)}_{(0,\ldots,0,\beta_{s+1},\ldots,\beta_{d})}}$
$\leq$ $\sup_{\beta\in\mathbb{N}_{0}^{d}}\log\frac{(Ah)^{\beta_{1}+\ldots+\beta_{s}}}{M^{(1/j)}_{(\beta_{1},\ldots,\beta_{s},0,\ldots 0)}}$
$+\sup_{\beta\in\mathbb{N}_{0}^{d}}\log\frac{(Ah)^{\beta_{s+1}+\ldots+\beta_{d}}\alpha_{s+1}^{\beta_{s+1}/2}\cdots\alpha_{d}^{\beta_{d}/2}}{M^{(1/j)}_{(0,\ldots,0,\beta_{s+1},\ldots,\beta_{d})}}$
(6.22) $\leq$ $\omega_{\bar{\mathbf{M}}^{(1/j)}}(Ah)+\omega_{\mathbf{M}^{(1/j)}}(j\bar{\alpha}^{1/2})$
for $\{\bar{M}^{(1/j)}_{\beta}\}_{\beta\in\mathbb{N}_{0}^{s}}$ defined by $\bar{M}^{(1/j)}_{\beta}:=M^{(1/j)}_{(\beta_{1},\ldots,\beta_{s},0,\ldots,0)}$, where we have chosen $j\geq Ah$, taking into account that the associate function is increasing on $(0,+\infty)^{d}$. Note also that if $j\geq j_{0}$ then $\mathbf{M}^{(1/j)}\leq\mathbf{M}^{(1/{j_{0}})}$ and hence $\omega_{\mathbf{M}^{(1/j)}}\geq\omega_{\mathbf{M}^{(1/{j_{0}})}}$. Since $\omega_{\bar{\mathbf{M}}^{(1/j)}}(Ah)$ is a new constant depending on $\ell$ ($A$ and $j$ depend on $h$ that depends on $\ell$), substituting (6.22) into (6.21) we have that
(6.23) $\begin{split}&\forall\ell\in\mathbb{N}\,\exists j\in\mathbb{N},C\geq 1\,\mbox{s.t.}\,\forall\mathbf{c}\in\Lambda_{(\mathcal{M})}\\ &\sup_{\alpha\in\mathbb{N}_{0}^{d}}|c_{\alpha}|e^{\omega_{\mathbf{N}^{(1/\ell)}}(\ell\alpha^{1/2})}\leq C\sup_{\alpha\in\mathbb{N}_{0}^{d}}|c_{\alpha}|e^{\omega_{\mathbf{M}^{(1/j)}}(j\bar{\alpha}^{1/2})}.\end{split}$
For $\beta\in\mathbb{N}_{0}^{d}$ we now consider $\mathbf{c}^{\beta}=(c_{\alpha}^{\beta})_{\alpha\in\mathbb{N}_{0}^{d}}$ defined by
$c_{\alpha}^{\beta}:=\delta_{\alpha\beta}=\begin{cases}1&\mbox{if}\ \alpha=\beta\cr 0&\mbox{if}\ \alpha\neq\beta.\end{cases}$
Clearly $\mathbf{c}^{\beta}\in\Lambda_{(\mathcal{M})}$ since
$\sup_{\alpha\in\mathbb{N}_{0}^{d}}|c_{\alpha}^{\beta}|e^{\omega_{\mathbf{M}^{(1/h)}}(h\alpha^{1/2})}=e^{\omega_{\mathbf{M}^{(1/h)}}(h\beta^{1/2})}<+\infty$
for every $\beta\in\mathbb{N}_{0}^{d}$ and $h\in\mathbb{N}$, by assumption (6.4) which implies that $\mathbf{M}^{(1/h)}$ satisfies (5.1) and hence $\omega_{\mathbf{M}^{(1/h)}}(t)<+\infty$ for all $t\in\mathbb{R}^{d}$.
We apply (6.23) to the sequences $\mathbf{c}^{\beta}$ and get that for all $\ell\in\mathbb{N}$ there exist $j\in\mathbb{N}$ and $C\geq 1$ such that for all $\beta\in\mathbb{N}_{0}^{d}$ we have
$e^{\omega_{\mathbf{N}^{(1/\ell)}}(\ell\beta^{1/2})}\leq Ce^{\omega_{\mathbf{M}^{(1/j)}}(j\bar{\beta}^{1/2})}$
or equivalently
(6.24) $\omega_{\mathbf{N}^{(1/\ell)}}(\ell\beta^{1/2})\leq\log C+\omega_{\mathbf{M}^{(1/j)}}(j\bar{\beta}^{1/2}).$
If $t\in(0,+\infty)^{d}$ then there exist $\beta\in\mathbb{N}_{0}^{d}$ with $\beta^{1/2}<t\leq(\beta+\mathbf{1})^{1/2}$ so that, from (6.24):
$\omega_{\mathbf{N}^{(1/\ell)}}(\ell t)\leq$ $\omega_{\mathbf{N}^{(1/\ell)}}(\ell(\beta+\mathbf{1})^{1/2})\leq\log C+\omega_{\mathbf{M}^{(1/j)}}(j(\overline{\beta+\mathbf{1}})^{1/2})$
$\leq$ $\log C+\omega_{\mathbf{M}^{(1/j)}}(2j\beta^{1/2})\leq\log C+\omega_{\mathbf{M}^{(1/j)}}(2jt)$
since the associate function is increasing on $(0,+\infty)^{d}$ and $(\overline{\beta+\mathbf{1}})^{1/2}\leq 2\beta^{1/2}$.
It finally follows from (5.6) and (5.7) that for all $\ell\in\mathbb{N}$ there exist $j\in\mathbb{N}$ and $C\geq 1$ such that
$N_{\alpha}^{(1/\ell)}\geq$ $\sup_{t\in(0,+\infty)^{d}}\frac{t^{\alpha}}{\exp\omega_{\mathbf{N}^{(1/\ell)}}(t)}=\sup_{s\in(0,+\infty)^{d}}\frac{(s\ell)^{\alpha}}{\exp\omega_{\mathbf{N}^{(1/\ell)}}(s\ell)}$
$\geq$ $\frac{\ell^{|\alpha|}}{C}\sup_{s\in(0,+\infty)^{d}}\frac{s^{\alpha}}{\exp\omega_{\mathbf{M}^{(1/j)}}(2js)}=\frac{\ell^{|\alpha|}}{C}\sup_{t\in(0,+\infty)^{d}}\frac{\left(\frac{t}{2j}\right)^{\alpha}}{\exp\omega_{\mathbf{M}^{(1/j)}}(t)}$
$=$ $\frac{1}{C}\left(\frac{\ell}{2j}\right)^{|\alpha|}M_{\alpha}^{(1/j)}$
and we have thus proved that $\mathcal{M}(\preceq)\mathcal{N}$, since the sequences $\mathbf{M}^{(1/j)}$ and $\mathbf{N}^{(1/\ell)}$ are normalized. ∎
Acknowledgments. Boiti and Oliaro were partially supported by the INdAM - GNAMPA Project 2023 “Analisi di Fourier e Analisi Tempo-Frequenza di Spazi Funzionali ed Operatori”. Boiti was partially supported by the Projects FAR 2020, FAR 2021, FAR 2022, FIRD 2022 and FAR 2023 (University of Ferrara). Jornet is partially supported by the project PID2020-119457GB-100 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. Schindl is supported by FWF (Austrian Science fund) project 10.55776/P33417.